# High-fidelity Metalens Imaging via Aberration Correcting Transformers

### v1

@@ 3.4 Spatial and Transposed Self-Attention Fusion
As aforementioned, Restormer [92] relies solely on transposed self-attention to generate an attention map, often resulting in weak spatial context encoding. To improve context modeling, a variety of subsequent studies [13, 16, 81] leveraged both SA and TA, by alternatively applying them as depicted in Fig. 5 (a). In particular, X-Restormer [13] added spatial self-attention to Restormer and proved the enhanced spatial context encoding capability.
Nevertheless, there remains room for improvement, as existing methods apply SA and TA uniformly across both encoder and decoder stages despite their distinct objectives. The encoder primarily focuses on capturing a global context, emphasizing the overall structure and relationships within the images. This is critical in image restoration tasks since understanding the broader contexts helps identify patterns and features that may be corrupted in the degraded images. In contrast, the decoder mainly aims to recover local details and textures essential for high-fidelity restoration.
Therefore, we propose a Spatial and Transposed self-Attention Fusion (STAF) module to further ameliorate image restoration. STAF implements SA and TA separately, not alternatively, and fuses SA and TA features with different weights in the encoder and decoder, as depicted in Fig. 5 (b). In detail, the outcome of SA and TA after feed-forward networks (denoted as $F_{SA}$, $F_{TA}\in\mathbb{R}^{H\times W\times C}$, respectively and $H,W,C$ are height, width, and channel of the input feature) are concatenated along channel dimension to obtain concatenated feature $F\in\mathbb{R}^{H\times W\times 2C}$. $F$ is then fed to a small CNN, consisting of two 1$\times$1 convolution layers with GELU activation [34], to mix features and halve channel dimension to $C$, and produce $A\in\mathbb{R}^{H\times W\times C}$, which is then treated differently in encoder and decoder. At the encoder side, we employ average pooling on spatial dimensions to efficiently reduce resolution while preserving crucial feature representations and apply the sigmoid function to $A$ to obtain weight matrix $A_{enc}\in\mathbb{R}^{1\times 1\times C}$. On the other hand, $W$ is averaged along the channel dimension to selectively emphasize the most relevant feature maps and apply the sigmoid function to produce weight matrix $A_{dec}\in\mathbb{R}^{H\times W\times 1}$ for the decoder. This allows the decoder to prioritize different spatial areas, for instance, regions with more aberration can receive higher attention to be corrected. In a nutshell, $A_{enc}$ and $A_{dec}$ can be computed as follows:
@@
where $\odot$ denotes the Hadamard product, and the index for each layer is omitted for brevity. The residual image obtained after passing through all STAF blocks is added to the $I_{dec}^{k_{med}}$, an image deconvolved with $W_{k_{med}}$ where $k_{med}$ is a median of $k_{m}$, to get the aberration-corrected image $I_{ac}$.
We adopted Overlapping Cross Attention (OCA), which partitions query matrix with non-overlapping windows and key and value matrices with overlapping windows, from HAT [15] for our SA module, following X-Restormer [13]. In addition, we employed the Gated-Deconv Feed-Forward network from Restormer [92] for our feed-forward network (FFN). However, any SA and FFN can be a valid option.
Fig. 6 shows the LAM [30] analysis which implies that STAF can restore the target region entailing more information, leading to better restoration quality. X-Restormer [13] also employs both SA and TA but aggregates the features in a simple manner, resulting in lacking information involvement and inferior aberration correction.
Figure 6: LAM [30] comparison over different models. Top: Input image and LAM analysis. Red marks in the images represent the pixels involved in restoring a target region, noted with a red box in the input image. Bottom: Restored results.
Figure 7: Qualitative results on the aberrated Open Image V7 [48, 7] dataset.
@@ 4.1 Experimental Settings
We trained MetaFormer for 300K iterations with a batch size of 8 on 4 RTX A6000 GPUs. We set the number of filters $M=3$ and noise penalizing scalers $k_{1}=1e-2,k_{2}=1e-3,k_{3}=3e-4$. We followed the overall experimental settings of Restormer [92] and X-Restormer [13]. Please refer to Restormer and X-Restormer for more details. We constructed synthetically aberrated datasets following Eq. 1 with metalens PSF from  Tseng et al. [80], where $n$ in Eq. 1 is a sum of two different per-pixel noises, Gaussian and Poisson noise.
We conducted experiments across various tasks: aberration correction on synthetic (Sec. 4.2) and real (Sec. 4.3) images, video aberration correction (Sec. 4.4), and 3D reconstruction with aberrated images (Tab. 2). Ablation studies are delivered in Sec. 4.6. Detailed experimental settings and more experimental results are described in the supplementary material.
4.2 Synthetic Image Aberration correction
We synthesized aberration on Open Image V7 [48, 7] dataset with PSF from Tseng et al. [80] and sampled 12K and 1K images for training and validation, respectively [^2] Training codes and pre-trained weights for Chakravarthula et al. [9] were not available, so we could only compare with Tseng et al. [80].. Tab. 1 and Fig. 7 describe the quantitative and qualitative results on the aberrated Open Image V7 dataset, respectively. Eboli et al. [23], a blind deconvolution approach, failed to compensate for the aberrations, and other baseline models, including Restormer [92], struggled with restoring fine details. Meanwhile, the proposed method could ameliorate Restormer with a noticeable margin and outperformed Tseng et al. [80] designed for metalens images.
@@ 4.2 Synthetic Image Aberration correction
Wiener deconvolution [86] 25.08 0.6433 0.5164
Eboli et al. [23] 16.45 0.3602 0.8929
DWDN [21] 25.77 0.7320 0.3333
Tseng et al. [80] 27.56 0.7612 0.3374
Restormer [92] 28.92 0.7719 0.3039
Ours 32.16 0.8159 0.2810
Table 1: Quantitative results on the aberrated Open Image V7 [48, 7] dataset. Its training dataset was not available, and the hyperparameters for training were not clearly disclosed at the time of writing. Thus, we used the pre-trained weights for the evaluation.
4.3 Real Image Aberration Correction
We fabricated a metalens with the PSF from Tseng et al. [80] as shown in Fig. 1 (Left) and captured images to construct a real image dataset. It consists of 220 images, where 200 images are involved in fine-tuning the pre-trained MetaFormer, and 20 images are kept for validation. Fig. 1 (Right) shows the qualitative results on the metalens-captured images. Even though MetaFormer is trained only with the synthetic data, it can correct the aberration of the real image after fine-tuning with a small real dataset, which eases the cost of constructing a large-scale real image dataset. Please refer to the supplementary material for details about metalens fabrication, image capture setup, model fine-tuning, and more experimental results.
@@ 4.4 Video Aberration Correction
In this section, we assessed our method for the video aberration correction task. We employed VRT [56] as a baseline video restoration model and trained it with synthetically aberrated DVD [76] dataset. We incorporated the MAFG into VRT optimization (VRT + MAFG) and trained VRT with the output of pre-trained MetaFormer (VRT w/ Ours) from Sec. 4.2.  Tab. 2 and Fig. 8 show that VRT struggled with correcting the aberration of the input video while simply applying the proposed MAFG to VRT (VRT + MAFG) could mitigate the severe aberration and restore the fine details. Furthermore, VRT trained with MetaFormer outputs (VRT w/ Ours) could enjoy both temporal consistency and clean representations, reaching the best restoration quality.
[…]
VRT [56] 16.38 0.4876 0.8058
VRT + MAFG 25.55 0.7795 0.2347
Restormer [92] 27.47 0.7463 0.2997
Ours 27.91 0.7798 0.2892
VRT w/ Ours 28.00 0.8331 0.2408
Table 2: Quantitative results on the aberrated DVD dataset [76].
@@
3D-GS [41] 15.34 0.3037 0.8706 14.85 0.2885 0.8641
3D-GS w/ Ours 24.07 0.6961 0.2732 22.25 0.6362 0.3259
3D-GS + Ours 24.37 0.7009 0.2867 22.40 0.6427 0.3113
Table 3: Quantitative results on the aberrated LLFF [64] and Tanks&Temples [44] dataset.
4.5 3D Reconstruction with Aberrated Images
We further extended our method to clean 3D reconstruction, training 3D Gaussian Splatting (3D-GS) [41] with the sets of multi-view aberrated images. We embedded MetaFormer into 3D-GS training pipeline as delineated in Fig. 9. Specifically, we simulated aberration during 3D-GS training using Eq. 1 to constrain 3D-GS learning clean representations. The output ($I_{ac}$) of the pre-trained MetaFormer, which is frozen during 3D-GS training, is involved in supervising the rendered image ($\mathcal{G}_{c}$) with a loss $\mathcal{L}_{ac}=\mathcal{L}_{GS}(\mathcal{G}_{c},I_{ac})$ where $\mathcal{L}_{GS}$ is a reconstruction loss function defined in 3D-GS. The aberration-simulated image $\mathcal{G}_{abr}$ is used to compute $\mathcal{L}_{abr}=\mathcal{L}_{GS}(\mathcal{G}_{abr},I_{abr})$, with the training image $I_{abr}$. We use a sum of these losses to guide 3D-GS learning clean features from the aberrated images.
Tab. 3 and Fig. 10 provide the quantitative and qualitative results on the 3D reconstruction task, evaluated on LLFF [64] and Tanks&Temples [44] dataset, where we used Truck and Train data in Tanks&Temples dataset following 3D-GS. While 3D-GS fails to render clean images, the proposed pipeline that combines 3D-GS and MetaFormer (3D-GS + MetaFormer) shows sound reconstruction performance, even outperforming 3D-GS optimized with the restored output $I_{ac}$ from MetaFormer (3D-GS w/ MetaFormer). This is because 3D-GS + MetaFormer can learn clean representation via $\mathcal{L}_{ac}$ and meet the multi-view consistency with $\mathcal{L}_{abr}$ simultaneously, resulting in better reconstruction quality.
Figure 9: Training pipeline for MetaFormer embedded 3D-GS.
@@
4.6 Ablation Studies
We conducted ablation studies of the proposed methods. Tab. 4 shows the comparison of MAFG and STAF module. Applying a single Wiener filter improved Restormer [92] by 1.80 dB in PSNR, but it elevated more when coupled with MAFG and STAF. Other ablation studies are delivered in the supplementary material.
[…]
28.92
30.72
32.00
32.16
Table 4: Ablation study on the proposed methods. SWF stands for applying a single Wiener filter to Restormer [92].
@@ 5 Conclusion
We present MetaFormer, an aberration correcting transformer tailored for metalens imaging. We harness a Vision Transformer (ViT) that has not been fully explored in metalens imaging. We devise Multiple Adaptive Filters Guidance (MAFG) to enrich the feature of the aberration correcting network. The proposed Spatial and Transposed self-Attention Fusion (STAF) module further boosts aberration correction by implementing SA and TA separately and fusing the outcomes differently in the encoder and decoder. The comprehensive experiments across diverse tasks covering correcting aberrated images, videos, and 3D reconstruction with aberrated images show the proposed method achieving state-of-the-art performance on various tasks. We also fabricate a metalens and restore the images captured in the wild, proving the practicality of MetaFormer.
@@ A.2 Optical Setup for Imaging
Appendix B Additional Experiments and Results
In this section, we provide more experimental results on each task that were not delivered in the main paper due to the page limit. We also attached videos of the 3D reconstruction and video aberration correction task results.
Evaluation Metrics. We assessed the proposed method with various evaluation metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [84], and Learned Perceptual Image Patch Similarity (LPIPS) [94].
B.1 Synthetic Image Aberration Correction
Fig. 14 shows extra qualitative results on the aberrated Open Image V7 [48, 7] dataset. The proposed method could mitigate severe aberrations and restore fine details. We also provide noise maps in Fig. 15 and Restormer [92] showed noises on textureless regions or failed to match contrast, where noise and low contrast are especially more prominent in metalens imaging. On the other hand, MetaFormer could alleviate noises and achieve the correct contrast, reaching better results on multiple evaluation metrics.
B.2 Real Image Aberration Correction
Model Fine-tuning Details. We fine-tuned our model, initially trained on synthetically aberrated images in Sec. 4.2 in the main paper, using real images captured with the fabricated metalens. The fine-tuning dataset includes 200 real images for training and 20 images for validation. Despite the limited dataset size, fine-tuning MetaFormer for 3,500 iterations could produce promising and appealing results.
Results. Fig. 16 presents the qualitative results of images captured with the fabricated metalens and restored images using the proposed method. The captured images show harsh degradations and aberrations but MetaFormer could successfully restore fine details and textures. Such outcomes can demonstrate the adaptability of the proposed method to real-world scenarios.
B.3 Video Aberration Correction
Fig. 17 shows additional qualitative results on the aberrated DVD [76] dataset. Similar to Fig. 8 in the main paper, the proposed method could effectively address aberrations in the video and produce aberration corrected representations.
B.4 3D Reconstruction with Aberrated Images
More Method Details. We used the combined loss function $\mathcal{L}_{train}$, consisting of $\mathcal{L}_{abr}$ and $\mathcal{L}_{ac}$ to train the proposed ‘3D-GS + MetaFormer’as following:
$\mathcal{L}_{train}=\mathcal{L}_{abr}+\lambda\mathcal{L}_{ac},$ (7)
where $\lambda$ is a hyperparameter controlling weight of $\mathcal{L}_{ac}$. We used the pre-trained MetaFormer trained with the aberrated Open Image V7 [48, 7] dataset from Sec. 4.2 in the main paper to get aberration corrected images, $I_{ac}$.
Experimental Settings. For the LLFF [64] and Mip-NeRF360 [6] datasets, we set the densification threshold $D$ of 3D Gaussian Splatting (3D-GS) model [41] to $42e-$ 4, the pruning threshold $P=5e-3$ and the loss weighting hyperparameter $\lambda=1.0$. We also inactivated periodical opacity reset for the LLFF dataset to preserve more points in the forward-facing dataset [50]. We set $D=4e-4$, $P=1e-2$, and $\lambda=2.0$ for the Tanks&Temples [44] dataset.
Results. We conducted experiments on the aberrated Mip-NeRF360 [6]. Tab. 5 and Fig. 18 show the quantitative and qualitative results, respectively. Training 3D-GS with the output of MetaFormer (3D-GS w/ Ours) produced aberration alleviated representations but it suffered from artifacts due to the lack of multi-view consistency. In contrast, extra multi-view consistency guidance via $\mathcal{L}_{abr}$ could aid in mitigating those artifacts from 3D inconsistency and yield more clear representations.

### v2

@@ 3.4 Spatial and Transposed Self-Attention Fusion
As aforementioned, a variety of previous studies adopted spatial self-attention (SA) [51, 14], transposed self-attention (TA) [93], or both [12, 15, 81] for image restoration. Leveraging both SA and TA, by alternatively applying them as depicted in Fig. 5 (a), can encode spatial context effectively.
Nevertheless, there remains room for improvement, as existing methods apply SA and TA uniformly across both encoder and decoder stages despite their distinct objectives. The encoder primarily focuses on capturing a global context, emphasizing the overall structure and relationships within the images, which helps identify patterns and features. In contrast, the decoder mainly aims to recover local details and textures essential for high-fidelity restoration.
Therefore, we propose a Spatial and Transposed self-Attention Fusion (STAF) module which implements SA and TA separately, not alternatively, and fuses SA and TA features with different weights in the encoder and decoder, as depicted in Fig. 5 (b). In detail, the outcome of SA and TA (denoted as $F_{SA}$, $F_{TA}\in\mathbb{R}^{H\times W\times C}$, respectively and $H,W,C$ are height, width, and channel of the input feature) are concatenated along channel dimension to obtain concatenated feature $F\in\mathbb{R}^{H\times W\times 2C}$. $F$ is then fed to a small CNN, consisting of two 1$\times$1 convolution layers with GELU activation [30], to mix features and halve channel dimension to $C$, and produce $A\in\mathbb{R}^{H\times W\times C}$, which is then treated differently in encoder and decoder. At the encoder side, we employ average pooling to $A$ on spatial dimensions to efficiently reduce resolution while preserving crucial feature representations and obtain weight matrix $A_{enc}\in\mathbb{R}^{1\times 1\times C}$. On the other hand, $A$ is averaged along the channel dimension to selectively emphasize the relevant feature maps and produce a weight matrix $A_{dec}\in\mathbb{R}^{H\times W\times 1}$ for the decoder. This allows the decoder to prioritize different spatial areas. In a nutshell, $A_{enc}$ and $A_{dec}$ can be computed as follows:
@@
where $\odot$ denotes the Hadamard product, and the index for each layer is omitted for brevity.
To further compensate for fine detail losses due to the harsh aberration, we enforce the queries of both attention modules in the decoder to focus more on high-frequency information, inspired by ESRT [57] that estimates high-frequency information of the feature using average pooling. In particular, we adopt this enhanced query $Q_{dec}^{b}$ in every other block of each decoder layer, as follows:
$Q_{dec}^{b}=\begin{cases}Q_{dec}^{\prime b},&\text{if }b=2n\\ Q_{dec}^{\prime b}-\texttt{AVGPool}(Q_{dec}^{\prime b}),&\text{if }b=2n+1,\\ \end{cases}$ (7)
where $b$ is a block index, $n\in\mathbb{Z}$, $Q_{dec}^{\prime b}$ is an original query from the previous block, and the index of the layer is again omitted for brevity. This enhanced query is used in the SA and TA blocks of the decoder to compute $F_{dec}$. Finally, the residual obtained after forwarding all STAF blocks is added to the $I_{dec}^{k_{med}}$, to get the corrected image $I_{ac}$.
Figure 6: Generated point spread functions (PSFs) and an image with spatially varying aberration derived from them.
Figure 7: Qualitative results on the Open Image V7 [44, 6] dataset with spatially varying aberration.
Figure 8: Qualitative results on the Open Image V7 [44, 6] dataset with the uniform aberration at field angle 20°.
Figure 9: Qualitative results on the images captured with our fabricated metalens.
@@
Non-uniform aberration Uniform aberration at $\theta=0$° Uniform aberration at $\theta=20$°
[…]
Wiener deconvolution [86] 25.54 0.5743 0.5228 27.06 0.6561 0.4458 26.05 0.6332 0.5030
Eboli et al. [21] 15.19 0.3784 0.8270 20.73 0.5359 0.6488 18.55 0.4480 0.7973
DeblurGANv2 [43] 24.08 0.6863 0.3233 25.27 0.7338 0.3032 21.79 0.6048 0.4810
DWDN [19] 26.40 0.7656 0.2854 29.38 0.8318 0.2459 25.68 0.7282 0.3267
Tseng et al. [80] 28.66 0.8045 0.2949 29.88 0.8373 0.2507 28.09 0.7841 0.3117
Restormer [93] 27.84 0.8091 0.2753 31.09 0.8827 0.1804 28.53 0.8020 0.2871
Restormer [93] + SWF 29.87 0.8289 0.2812 33.58 0.8804 0.2003 28.95 0.7908 0.3262
Ours 34.29 0.8760 0.2052 35.06 0.8961 0.1763 32.32 0.8409 0.2542
Table 1: Quantitative results Open Image V7 [44, 6] dataset with spatially varying and invariant aberration, where $\theta$ denotes a field angle.
@@ 4.1 Experimental Settings
We divided FOV into $9\times 9$ grid and generated 81 PSFs using metalens phase profile from Tseng et al. [80] to represent spatially varying aberration of metalens imaging as shown in Fig. 6, following Eq. 1. We set the number of filters $M=13$ and selected 13 logarithmically spaced values in the range $[1e-5,1e-2]$ for noise penalizing scalers $k_{m}$.
We conducted experiments across various tasks: image aberration correction (Fig. 10), restoring images captured with the fabricated metalens (Sec. 4.3), video aberration correction (Sec. 4.4), and 3D reconstruction with aberrated images (Sec. 4.5). Analysis on the method, detailed experimental settings, and more experimental results are described in the supplementary material.
4.2 Image Aberration Correction
We adopted Open Image V7 [44, 6] dataset with spatially varying aberration for image aberration correction task, and sampled 120K and 1K images for training and validation, respectively. We constructed two different types of test sets, one with spatially varying aberrations from 81 different PSFs, and the other with spatially invariant aberration, where we selected 5 PSFs at different field angles (0°, 5°, 10°, 15°, and 20°), each corresponding to a distinct uniform aberration.
Tab. 1, Fig. 7 and Fig. 8 provide the results on the spatially varying and invariant aberration test sets [^1] The codes and pre-trained weights for Chakravarthula et al. [8] were not publicly available, so we could only compare with Tseng et al. [80]., where Restormer+SWF is a non-blindly trained Restormer using a single Wiener filter. As shown in input images in Fig. 8, an image degraded by the PSF at a large field angle suffers more from aberration, which corresponds to the outer parts of the image with spatially varying aberration in Fig. 7.
Spatially Varying Aberration Tseng et al. [80] leverages the PSF to address aberrations but fails to preserve fine details and contrast, highlighting the limitations of CNNs for metalens imaging. Restormer [93] demonstrates the capability of ViTs in metalens imaging by restoring the mildly aberrated region (blue box in Fig. 7) without PSF guidance. However, it fails to correct pronounced aberrations in off-axis areas (red box in Fig. 7). Non-blindly training Restormer with a single Wiener filter still struggles with recovering fine details. In contrast, our approach effectively corrects aberrations across both on- and off-axis regions, achieving state-of-the-art results with a noticeable margin.
Spatially Invariant Aberration Tseng et al. [80] models spatially varying aberration by utilizing multiple patches with different degrees of aberration. However, since each patch assumes a single type of aberration (i.e., uniform aberration), seam-like artifacts may emerge when applied to the image where diverse aberrations coexist within a single frame, which is common in real scenarios. To ensure a fair comparison, we further conducted experiments under uniform aberration conditions. As demonstrated in Tab. 1 and Fig. 8, our method effectively models spatially varying aberrations, achieving robust performance even in cases where severe aberrations (at field angle 20°) appear uniformly, outperforming Tseng et al. [80].
@@ 4.2 Image Aberration Correction
VRT [52] 23.23 0.6906 0.3921
VRT w/ Ours 28.89 0.8602 0.2102
Table 2: Quantitative results on the aberrated DVD dataset [75].
4.3 Images from the Fabricated Metalens
We fabricated a metalens with the PSF from Tseng et al. [80] as shown in Fig. 9 (Left) and captured images to construct a dataset. It consists of 220 images, where 200 images are involved in fine-tuning the pre-trained model from Fig. 10, and 20 images are kept for validation. Since ViT requires a larger volume of training data, large-scale dataset construction is inevitable, which is labor-intensive. Instead, we trained a model with numerous data simulated with the same phase profile of the manufactured metalens, and then fine-tuned it with the small set of captured images, easing the cost of collecting data. We fine-tuned our model, Restormer [93] and non-blind Restormer under the same experimental setting. As shown in Fig. 9, both Restormer and non-blind Restormer produce images in low contrast while the proposed method yields images with better contrast, leading to more visually appealing quality.
@@ 4.4 Video Aberration Correction
In this section, we assessed our method for the video aberration correction task. We employed VRT [52] as a baseline video restoration model and trained it on the DVD [75] dataset with spatially varying aberration. We trained VRT with the output of pre-trained model (VRT w/ Ours) from Fig. 10. Tab. 2 and Fig. 10 show that VRT struggled with correcting the aberration of the input video while VRT w/ Ours could enjoy both temporal consistency and clean representations.
4.5 3D Reconstruction with Aberrated Images
@@
3D-GS [37] 17.43 0.4083 0.7428 16.81 0.3875 0.7263
3D-GS w/ Ours 25.29 0.7513 0.2138 23.17 0.6852 0.2639
3D-GS + Ours 25.49 0.7548 0.2149 23.30 0.6905 0.2632
Table 3: Quantitative results on the LLFF [60] and Tanks&Temples [40] dataset with spatially varying aberration.
Figure 11: Training pipeline for 3D-GS using the proposed ViT.
@@
We further extended our method to clean 3D reconstruction, training 3D Gaussian Splatting (3D-GS) [37] with the sets of multi-view aberrated images. We embedded our method into 3D-GS training pipeline as delineated in Fig. 11. Specifically, we simulated aberration during 3D-GS training using Eq. 1 to constrain 3D-GS learning clean representations. The output ($I_{ac}$) of our pre-trained ViT, which is frozen during 3D-GS training, is involved in supervising the rendered image ($\mathcal{G}_{c}$) with a loss $\mathcal{L}_{ac}=\mathcal{L}_{GS}(\mathcal{G}_{c},I_{ac})$ where $\mathcal{L}_{GS}$ is a reconstruction loss function defined in 3D-GS. The aberration-simulated image $\mathcal{G}_{abr}$ is used to compute a loss $\mathcal{L}_{abr}=\mathcal{L}_{GS}(\mathcal{G}_{abr},I_{abr})$, with the training image $I_{abr}$. We use a sum of these losses, $\mathcal{L}_{train}=\mathcal{L}_{abr}+\lambda\mathcal{L}_{ac}$, to guide 3D-GS learning clean features from the aberrated images, where $\lambda$ is a hyperparameter.
Tab. 3 and provide the results on the 3D reconstruction task, evaluated on LLFF [60] and Tanks&Temples [40] dataset with spatially variant aberration where we used only Truck and Train scenes in Tanks&Temples dataset for evaluation, following the experimental setting in 3D-GS. While 3D-GS fails to render clean images, the proposed pipeline that combines 3D-GS and our image restoration model (3D-GS + Ours) shows sound reconstruction performance, even outperforming 3D-GS optimized with the restored output $I_{ac}$ from the proposed image correction model (3D-GS w/ Ours). This is because 3D-GS + Ours can learn clean representation via $\mathcal{L}_{ac}$ and multi-view consistency with $\mathcal{L}_{abr}$ simultaneously, resulting in better reconstruction quality. Fig. 12 shows qualitative results, where the proposed method can produce clean images in novel views, while vanilla 3D-GS still suffers from severe aberration.
@@ 5 Conclusion
We present an aberration correcting transformer tailored for metalens imaging. We harness a Vision Transformer (ViT) that has not been fully explored in metalens imaging. We devise Multiple Adaptive Filters Guidance (MAFG) that can better model spatially varying aberration of metalens images and Spatial and Transposed self-Attention Fusion (STAF) which exploits features from SA and TA considering the roles of the encoder and decoder. The comprehensive experiments show the proposed method achieving state-of-the-art performance across diverse tasks. We also fabricate a metalens and effectively restore the captured images, proving the practicality of the proposed method.
@@ A.2 Optical Setup for Imaging
Appendix B Analysis and Discussion
In this section, we provide a comprehensive analysis to demonstrate the effectiveness of the proposed methods. All models are trained with a quarter of the original batch size.
Effect of MAFG. To validate the effectiveness of MAFG, we conducted ablation studies on each module of MAFG as shown in Tab. 4, which compares MAFG and other two baselines. The first baseline (MAFG w/o CA) does not apply cross-attention to the deconvolved images but simply concatenates them and feeds the concatenated features to the restoration network, which drops PSNR by 0.28 dB. We assume this is because cross-attention can reweight the deconvolved representations considering the degree of aberrations, proving it is proper to handle metalens images where aberration varies across the image. The other baseline (MAFG w/o CI) does not consider channel intensity when designing Wiener filters. Noise is signal-dependent (e.g. bright regions have more noise variance) and requires different noise regularization on different intensities, but this approach cannot reflect such property. It undermines perceptual quality, degrading LPIPS by 0.0038. We believe that a more sophisticated approach to incorporating image information into the filters can further improve the restoration quality. Lastly, Tab. 5 compares the number of filters $M$ involved in MAFG. The more filters we employed, the aberration correction performance was improved as it could enrich the input of the restoration network more and reweight the features more adaptively. Thus, we use 13 different filters considering the memory.